A commuter backs her car out of her garage with an acceleration of 2m/s^2. How long does it take her to reach a speed of 2m/2? Round answer in two decimal places


1
Expert's answer
2021-02-15T17:42:16-0500
"v=v_0+at,""v_0=0, t=\dfrac{v}{a}=\dfrac{2\ \dfrac{m}{s}}{2\ \dfrac{m}{s^2}}=1.0\ s."